OpenStudy (abbyrose0105):

Simplify 10 over quantity of 8 minus 5 i.

OpenStudy (kirbykirby):

Multiply by the conjugate \[ \frac{10}{8-5i}=\frac{10}{8-5i}\frac{8+5i}{8+5i}\]

OpenStudy (abbyrose0105):

Alright after that then what?

OpenStudy (kirbykirby):

multiply out the numerator and denominator

OpenStudy (abbyrose0105):

How do I do that?

OpenStudy (kirbykirby):

just like with regular numbers... \[ \frac{10(8+5i)}{(8-5i)(8+5i)}\]

OpenStudy (abbyrose0105):

Would you mind breaking it down a little more? I've had 4 hours of sleep in the last two days haha I'm a little slow today

OpenStudy (kirbykirby):

\[\frac{10(8+5i)}{(8-5i)(8+5i)}=\frac{10(8)+10(5i)}{8(8)+8(5i)-5i(8)-5i(5i)}\]

OpenStudy (abbyrose0105):

so the answer is quantity of 80 plus 50 i over 89 ??

OpenStudy (kirbykirby):

yup

OpenStudy (abbyrose0105):

awesome thanks

OpenStudy (kirbykirby):

yw :)